According to the table at right, the four-firm concentration ratio of this industry is O A. 69.2 peroent. OB. 35.1 percent. O C. 67.5 percent. O D. 66.7 percent. Annual Sales $1,000 Firm 900 120 75 40 800 1,200 1,050 90 75 600
Answer #1

Four firm concentration ratio = Sum of annual sales of four firms which earns highest / Total sales

= (1200 + 1050 + 1000 + 900)/(1000 + 900 + 120 + 75 + 50 + 40 + 800 + 1200 + 1050 + 90 + 75 + 600)

= 4150/6000 = 69.2 %

Answer is A) 69.2 percent